OpenStudy (umarsback):

Write an equation of a line in slope intercept form that is parallel to the line x + 4y = 6 and passes through (-8, 5).

OpenStudy (anonymous):

man. i can't do all your vvork lol

OpenStudy (umarsback):

im sorry, theres only 2 more questions i need to do :/

OpenStudy (anonymous):

Here is the explaination http://cl.ly/image/41021W1f2j1B

OpenStudy (anonymous):

You're going to need help on the point slope form. its actually very very easy.

OpenStudy (anonymous):

well do you know what slope intercept for is first of all

OpenStudy (anonymous):

find the slope of equatin x+4y=6 for that you should fine the coefficient of x then wrt equation with taht slope y-5=m(x+8) now just find m :)

OpenStudy (anonymous):

first know what slope intercept for is and make sure you understand CLEARLY what slope means